OpenStudy (kawii2004):

In which measurement situations would the most appropriate unit of measurement be inches? Choose exactly two answers that are correct. A. width of the Mississippi River B. diameter of a pizza C. height of a skyscraper D. length of a pencil

OpenStudy (studio-games):

D

OpenStudy (studio-games):

Mind medaling and fanning me please?

OpenStudy (studio-games):

Oh and B

OpenStudy (studio-games):

D and B

OpenStudy (anonymous):

no its just D. nice try though

OpenStudy (studio-games):

Choose exactly TWO answers that are correct... READ

OpenStudy (anonymous):

you aint gotta be a wingspan about it

OpenStudy (kawii2004):

james i belive u

OpenStudy (studio-games):

...

OpenStudy (anonymous):

well it is both

OpenStudy (kawii2004):

whats the awnser?

OpenStudy (studio-games):

D and B

OpenStudy (studio-games):

Pizza and pencil.

OpenStudy (kawii2004):

A rectangle has a length of 10 cm and a perimeter of 46 cm. Which equation could be used to find the width of the rectangle? A. 10w = 46 B. w + 10 = 46 C. 2w + 20 = 46 D. 2w + 10 = 46

OpenStudy (studio-games):

C

OpenStudy (anonymous):

it is C.

OpenStudy (studio-games):

Yeah its C